tavon has a gift card for $125 that loses $3 each 30-day period it is not used. He has another gift card for $105 that loses $2.50 for each 30-day period it is not used.


1
Expert's answer
2021-09-14T06:05:04-0400

"125-3\times x=105-2.5\times x\\125-105=3\times x- 2.5\times x\\20=0.5\times x\\x=40"


so the value of each gift card will be

"125-3\times x=105-2.5\times x"

now putting the value of x in above equation

"125-3\times 40=105-2.5\times 40\\125-105=105-100\\5=5"

so the value of each gift card will be $5.